# From Competition to Centralization: The Oligopoly in Ethereum Block Building Auctions

### v1

@@ 4.1. $\alpha$-Rank
In the first game, where players share uniform latency and orderflow access distribution, we begin with a small $\alpha$ value and increase it exponentially (i.e., sweep), following the approach suggested in Omidshafiei et al. (2019). However, as the size of the game grows, this sweeping becomes computationally challenging. Thus, we use a method from OpenSpiel Lanctot et al. (2019) to estimate a lower-bound value of $\alpha$ [^3] https://github.com/google-deepmind/open_spiel/pull/403.
5. Results
In this section, we present the results of the three empirical games. We examine the equilibria of these games, focusing on the strategic incentives of builders under different conditions. We analyze how advantages in latency and orderflow access influence builders’ incentives, and assess the overall state of the MEV-Boost auction. We discuss how our findings compare to those in Wu et al. (2024b) in Section 6.
5.1. Symmetric game
@@
Figure 1. Mass of all strategy profiles in the stationary distribution when sweeping $\alpha$. The legend shows the final ranking of strategy profiles as computed by $\alpha-$Rank.
[…]
@@
5.2. Latency impact
In the role-symmetric games, all players maintain the same prior distribution of private orderflow access, but experience two different types of latencies (low and high). The key factor affecting builders’ strategy choices is the latency gap between low- and high-latency groups. To analyze this, we consider scenarios where the latency difference between the two groups ranges from 0ms (previous symmetric scenario) to 200ms, increasing in 10ms increments. The latency of low-latency players remains fixed at 10ms, while the latency of high-latency players starts at 10ms and increases by 10ms in each subsequent scenario. Each scenario is treated as an independent game. We analyze the equilibria of these games to assess the impact of latency differences.
@@ 5.2. Latency impact
Figure 2. Average usage of aggressive and the other two (conservative+moderate) strategies by low-latency players (left) and high-latency players (right) across all profiles under varying latency differences as computed by $\alpha$-Rank.
[…]
Average usage of strategies.
@@
We further compute the Herfindahl-Hirschman Index (HHI) to measure the level of market centralization at equilibrium Leonardos et al. (2020). The win rate of each individual player is assumed to be evenly distributed within their respective groups. For each group, the individual win rates are:
@@
Figure 3. Overall win rates (left) of low- and high-latency players and HHI (right) under varying latency differences.
[…]
Latency winrate
@@
Figure 4. The overall auction efficiencies under varying latency differences.
[…]
Latency efficiencies.
Although further increasing the latency difference between the low- and high-latency players can affect the players’ incentive, particularly for low-latency players, and eventually shift the equilibrium, empirical studies Öz et al. (2024) have shown that, among the top ten builders, the average latency difference between the five faster builders and five slower builders is typically around 200ms. This indicates that the setup of our games provides practical insights into the real market.
5.3. Orderflow access impact
@@
Figure 5. Average usage of each meta strategy by high-orderflow players (left) and low-orderflow players (right) across all profiles under varying orderflow access probability differences as computed by $\alpha$-Rank.
[…]
Various orderflow.
Figure 6. Overall win rates (left) of low- and high-orderflow players and HHI (right) under varying orderflow access probability differences.
[…]
Various latencies - win rates.
@@
To counter this trend, smaller builders, with limited access to private orderflow, are often forced to bid aggressively, making minimal profits or even subsidizing bids to maintain their presence in the market Öz et al. (2024), which is financially unsustainable in the long term. This observed dynamic raises concerns about the long-term health of the market, as it discourages new entrants and entrenches the position of established builders, contributing to further centralization of the Ethereum block building market Yang et al. (2024).
@@
Figure 7. The overall auction efficiencies under varying orderflow access probabilities of high-orderflow builders.
[…]
Auction efficiency - orderflow.
6. Discussion
Our results show that under idealized conditions—where builders have equal latency and access to orderflow—the block building market is competitive and decentralized, and the MEV-Boost auction is efficient. All builders bid aggressively and achieve equal market share, and a significant portion of MEV is captured by proposers. However, achieving such idealized conditions under the current PBS and orderflow landscape is particularly challenging. The primary challenge arises from the unequal distribution of orderflow caused by vertical integration across the MEV supply chain. Vertical integration allows builders to access more frequent and valuable orderflow, significantly improving their performance in auctions. As indicated by our results, both advantages in latency and orderflow access incentivize dominant builders to bid conservatively while maintaining high profit margins, contributing to market centralization and auction inefficiency.
For MEV searchers and orderflow providers, vertical integration is primarily motivated by the lack of a trustless and private mechanism for distributing orderflow. In the absence of such a system, orderflow providers rely on exclusive deals with builders in a trusted and permissioned manner to protect their MEV from being stolen. It is common for orderflow providers to send their orderflow exclusively to large, dominant builders, as this maximizes the chances of their transactions landing on-chain while safeguarding their MEV. This practice creates an imbalanced distribution, where large builders capture the majority of orderflow, leaving smaller builders with limited access and diminishing their competitiveness in the auction market.
Efforts to address this issue include Orderflow Auctions, like MEV-Share Flashbots (2024c) and MEVBlocker Blocker (2024), which aim to provide fairer distribution mechanisms while protecting MEV for providers. However, these systems are still operated by trusted parties. A promising solution under exploration is BuilderNet Flashbots (2024a), which is a globally distributed block building network that leverages Trusted Execution Environments (TEEs) to enable private transaction sharing, decentralized block building by multiple operaters (builders), and MEV redistribution. However, since the project is in its early stages, its significance in the market structure is currently unclear.
6.1. Impact of strategy space
In this paper, we revisit the MEV-Boost auction games with similar settings to those studied in Wu et al. (2024b), but with a richer strategy space. By introducing a set of meta-strategies that cover the entire strategy space of builders, we are able to test the robustness of the results obtained in Wu et al. (2024b) under the limited strategy space of truthful, adaptive, and last-minute strategies. In the setting of Wu et al. (2024b), all builders converged to an adaptive bidding strategy, marginally outbidding each other and achieving significant profits, which led to low auction efficiency. Even in scenarios with orderflow disparities, their results showed little incentive for builders to deviate from this strategy.
Our findings paint a different picture showcasing the impact of strategy choices in the MEV-Boost auction outcomes. We demonstrate that under a richer strategy space builders have different incentives, changing both the competitive dynamics and the auction outcomes. These findings align with the current real-world data and offer novel insights into the underlying key dynamics and builders’ incentives that drive the outcomes in current MEV-Boost auctions.
6.2. Limitations
To ensure tractable equilibria with available computational resources, we typically modeled MEV-Boost auction games as one-stage auctions with a generalized strategy space that contains 3 meta strategies, and limited asymmetries between players. While the real-world MEV-Boost auction dynamics are more complex, this approach is rich enough to recover emerging behaviors and current trends in actual MEV-Boost auctions and provides meaningful insights into key factors such as builder incentives, auction efficiency, and market (de)centralization. A natural extension would be to partition the three bidding intervals into more intervals to expand the set of meta strategies; however, this approach did not alter our core findings and significantly increased analytical complexity. For instance, in the first experiment under idealized scenarios, one would imagine that builders remained incentivized to adopt the most aggressive strategy available among the intervals.
For future work, more sophisticated models can be developed to capture these intricate dynamics, such as incorporating consecutive auctions and builders’ bid-subsidizing strategies. Additionally, as we currently fix players’ strategies, future studies could implement reinforcement learning or other adaptive methods to enable strategy exploration, allowing for the simulation of best-response behavior and evolving strategies in MEV-Boost auctions.
7. Conclusion
@@
In this paper, we examined builders’ strategic bidding incentives in Ethereum block building auctions. Our results show that under idealized market conditions—where builders have comparable network latency and equal access to MEV opportunities—builders are incentivized to bid aggressively, resulting into a competitive and decentralized market with efficient auctions. However, when builders have unequal resources, as is the case in practice, the market tends to centralize. Well-resourced builders, especially those with superior access to private orderflow, are favored to benefit from economy of scale with reinforcing feedback loops. This creates an oligopolistic dynamic, leading to reduced proposer revenue and lower auction efficiency. Our results highlight the importance of fair MEV distribution among builders and underscore the necessity of enhancing decentralization in the Ethereum block building market.

### v2

@@ 4.1 $\alpha$-Rank
In the first game, where players share uniform latency and orderflow access distribution, we begin with a small $\alpha$ value and increase it exponentially (i.e., sweep), following the approach suggested in [21]. However, as the size of the game grows, this sweeping becomes computationally challenging. Thus, we use a method from OpenSpiel [17] to estimate a lower-bound value of $\alpha$. [^2] https://github.com/google-deepmind/open_spiel/pull/403.
5 Results
In this section, we present the results of the three games. We examine the equilibria of these games, focusing on the strategic incentives of builders under different conditions. We analyze how advantages in latency and orderflow access influence builders’ incentives, and assess the overall state of the MEV-Boost auction. We discuss how our findings compare to those in [35] in Section˜6.
5.1 Symmetric game
@@
Figure 1: Mass of all strategy profiles in the stationary distribution when sweeping $\alpha$. The legend shows the final ranking of strategy profiles as computed by $\alpha$-Rank.
@@
5.2 Latency impact
In the first role-symmetric game, all players maintain the same prior distribution of private orderflow access, but experience two different types of latencies (low and high). The key factor affecting builders’ strategy choices is the latency gap between low- and high-latency groups. To analyze this, we consider scenarios where the latency difference between the two groups ranges from 0ms (previous symmetric scenario) to 200ms, increasing in 10ms increments. The latency of low-latency players remains fixed at 10ms, while the latency of high-latency players starts at 10ms and increases by 10ms in each subsequent scenario. Each scenario is treated as an independent game. We analyze the equilibria of these games to assess the impact of latency differences.
@@ 5.2 Latency impact
Figure 2: Average usage of aggressive and the other two (conservative+moderate) strategies by low-latency players (left) and high-latency players (right) across all profiles under varying latency differences as computed by $\alpha$-Rank.
@@
We further compute the Herfindahl-Hirschman Index (HHI) to measure the level of market centralization at equilibrium [18]. The win rate of each individual player is assumed to be evenly distributed within their respective groups. For each group, the individual win rates are:
@@
Figure 3: Overall win rates (left) of low- and high-latency players and HHI (right) under varying latency differences.
@@
Figure 4: The overall auction efficiencies under varying latency differences.
Although further increasing the latency difference between the low- and high-latency players can affect the players’ incentive, particularly for low-latency players, and eventually shift the equilibrium, empirical studies [22] have shown that, among the top ten builders, the average latency difference between the five faster builders and five slower builders is typically around 200ms. This indicates that the setup of our games provides practical insights into the real market.
5.3 Orderflow access impact
@@
Figure 5: Average usage of each meta strategy by high-orderflow players (left) and low-orderflow players (right) across all profiles under varying orderflow access probability differences as computed by $\alpha$-Rank.
Figure 6: Overall win rates (left) of low- and high-orderflow players and HHI (right) under varying orderflow access probability differences.
@@
To counter this trend, smaller builders, with limited access to private orderflow, are often forced to bid aggressively, making minimal profits or even subsidizing bids to maintain their presence in the market [22], which is financially unsustainable in the long term. This observed dynamic raises concerns about the long-term health of the market, as it discourages new entrants and entrenches the position of established builders, contributing to further centralization of the Ethereum block building market [38].
@@
Figure 7: The overall auction efficiencies under varying orderflow access probabilities of high-orderflow builders.
6 Discussion
Our results show that under idealized conditions –where all builders share equal latency and orderflow access– the block building market remains competitive and decentralized, with an efficient MEV-Boost auction. However, in practice, exclusive orderflow deals arise largely because orderflow providers have no trustless, private way to distribute flow; they form exclusive agreements with builders to protect their MEV and secure priority block inclusion. Consequently, large builders receive preferential orderflow, while smaller builders face limited access and struggle to win. Vertically integrating with MEV searchers and relays further grants certain builders more frequent and valuable orderflow, enabling them to bid conservatively yet maintain high margins. These advantages drive centralization and reduce auction efficiency.
Recent proposals like MEV-Share [12], MEVBlocker [2], and BuilderNet [10] aim to provide a fairer and private orderflow distribution. However, these systems are either relying on trusted entities or remain in early development, leaving their long-term impact on market structure unclear.
Robustness of results.
Unlike [35], where their findings are inherently constrained by their restricted strategy space and not observed in real-world dynamics [22, 38], the main advantage of the meta-game framework is that it allows us to explore the strategy space in a more nuanced way while keeping the analysis tractable. In our case, this is achieved through the analysis of three meta-strategies that partition the entire strategy space. Despite the complexity of real-world MEV-Boost auctions, this model is rich enough to recover emerging behaviors and provides meaningful insights into builder incentives, auction efficiency, and market (de)centralization. Importantly, our current set of results is robust to the selected partition of the strategy space; our tests (not presented here) suggest that finer partitions do not qualitatively change the results while significantly increasing the analytical complexity.
7 Conclusion
@@
In this paper, we examined builders’ strategic bidding incentives in Ethereum block building auctions through empirical game theoretic analysis within the meta-game framework. Our results show that under idealized market conditions builders are incentivized to bid aggressively, resulting into a competitive and decentralized market with efficient auctions. However, when builders have unequal resources, as is the case in practice, the market tends to centralize. Well-resourced builders, especially those with superior access to private orderflow, are favored to benefit from economy of scale. This creates an oligopolistic dynamic, leading to reduced proposer revenue and lower auction efficiency. Our results highlight the importance of fair MEV distribution among builders and underscore the necessity of enhancing decentralization in the Ethereum block building market.
From a multi-agent system perspective, our results show how the meta game framework, allowing to account for an infinite number of strategies, qualitatively changes the findings of previous literature [35] in the domain: In MEV boost auctions, builders do not collude, but instead compete to create an oligopoly.
{ack}
We thank Julian Ma of Ethereum Foundation and one AFT reviewer for their insightful comments on the early version of this work.